OpenStudy (anonymous):

how many gallons of 60% antifreeze solution must be mixed with 60 gallons of 20% antifreeze to get a mixture that is 50% antifreeze

OpenStudy (anonymous):

60/100.x + 20/100.60=(x+60).50/100 3/5.x + 12 = 1/2.x +30 3/5.x - 1/2.x = 30-12 1/10.x = 18 x = 180 gallons

OpenStudy (anonymous):

60x + 20* (60) = 50* (x+60) 60x + 1200 = 50x + 3000 10x= 3000- 1200 = 1800/10 = 180 which means fatonbina was right....lolzzz :);)